OpenStudy (anonymous):

Let h = vo^2/4.9 sin(theta) cos(theta) model the horiontal distance in meters traveled by a projectile. If the initial velocity is 36 meters/ second, which equation would you use to find the angle needed to travel 100 meters? A. 100sin(2theta)= 100 B. 132.12sin(2theta)=100 c. 7.36sin(2Theta) = 100 D. 264.49sin(2theta) = 100

OpenStudy (anonymous):

\[\sin \theta \cos \theta=1/2\sin 2 \theta\] \[h=\frac{ v_0 \sin 2\theta }{ 4,9\times 2 }\]

OpenStudy (anonymous):

calculate and \[36/9.8\]see wich one mathces

OpenStudy (anonymous):

whatare you talking about???

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

hmm, after some thoght, does travel 100 mean "land at 100"? which is when h=0

OpenStudy (anonymous):

wait. ._. what? .... yeah ..that totally... ..great T.T

OpenStudy (anonymous):

i think its h= 100?

OpenStudy (amistre64):

ugh, internet is being mean to me ...

OpenStudy (anonymous):

ahhh

OpenStudy (amistre64):

the general formula for finding range of a projectile is: \[\frac{ (v_0)^2 \sin (2\theta) }{ g}=range\]

OpenStudy (amistre64):

so 36^2/9.8 sin(2 theta) = 100 what is 36^2/9.8 ??

OpenStudy (anonymous):

132.24

OpenStudy (amistre64):

and the closest option to that would start with 132.12 the difference is prolly due to a roundoff error on their part

OpenStudy (anonymous):

mmk